. The probability that a married man watches a certain television show is 0.4 and the probability that a

married woman watches the show is 0.5. The probability that a man watches the show, given that his

wife does is 0.7. Find the probability that

(a) a married couple watches the show;

(b) a wife watches the show given that her husband does;

(c) at least 1 person of a married couple will watch the show.


1
Expert's answer
2021-12-17T08:35:20-0500

Solution;

Given;

"P_M=0.4"

"P_W=0.5"

"P_{M|W}=0.7"

(a)

"P_{MW}=P_{M|W}\u00d7P_W=0.7\u00d70.5=0.35"

"(b)"

"P_{W|M}=\frac{P_{MW}}{P_M}=\frac{0.35}{0.4}=0.8750"

"(c)"

"P_{MorW}=P_M+P_W-P_{MW}=0.4+0.5-0.35=0.55"